A flat circular plate of copper has a radius of 0.243m. It's mass is 62kg. Friend the thickness of the plate. [the density of the copper is known to be 8.93g/cm³]

1
Expert's answer
2022-02-09T06:28:20-0500

The thickness is


"t=\dfrac{m}{\rho \pi r^2}=0.0374\text{ m}."